# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ D.3.1. Critical point in $I_{\alpha}^{\prime}$
(E.34) $\begin{split}&|\mathcal{Q}_{\epsilon}(u)|\leq c\epsilon\ln(1/\epsilon)(1+u^{2})^{-1},\ u\in\mathbb{R}.\end{split}$
For large $|u|$, (E.34) is proven integrating by parts in (E.33). For $|u|$ in a bounded set, we argue similarly to (E.12), (E.13).
Using (E.32) and (E.33), (E.34) with $h_{\epsilon}(u)=h(u;\theta,\epsilon)-\mu h_{0}(u)$ we get from (E.30)
(E.35) $\begin{split}&|G(\check{t},\check{q};\theta,\epsilon)-\mu G_{0}(\check{t},\check{q})|\leq c|J_{\epsilon}|+O(\epsilon^{1/2}),\\ &J_{\epsilon}:=\sum_{j_{2}}\mathcal{Q}_{\epsilon}(\check{t}+\check{q}-j_{2})(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon\ln(1/\epsilon)).\end{split}$
@@
We evaluate the integrals in (7.17) similarly to [25, eqs. (3.16)–(3.20)]. Begin by considering
@@
Substitute $h_{0}$ expressed in terms of its Fourier transform, $\tilde{h}_{0}=\mathcal{F}h$, and use the identity [14, p. 360, eq. 20]:
(E.37) $\mathcal{F}(x_{+}^{a-1})=e^{ai\pi/2}\Gamma(a)(\lambda+i0)^{-a},$
@@
Substitution of (E.39) back into (7.17) gives (7.18).
@@
Integrating by parts in (8.6) with respect to $\lambda$ twice and using a partition of unity we can assume that $I=(-\delta,\delta)$ for some small $0<\delta\ll 1$, and $D(\lambda,\theta,\epsilon)\in C_{0}^{\infty}(I)$ in $\theta$. On $I$ we can change variable $\theta\to t$ to get
@@
and $\tilde{D}_{1}(\lambda,\mu,\epsilon)$ is the Fourier transform of $D(\lambda,\theta,\epsilon)$ with respect to $\theta$. Thus $\tilde{D}_{1}(\lambda,\mu,\epsilon)$ decays rapidly as $\mu\to\infty$, i.e. for any $N\in\mathbb{N}_{0}$ there exists $c_{N}$ such that
@@
Changing variable $\lambda\to r\lambda$ proves the top cases in (8.7).
All the other assertions of the lemma can be proven in a similar fashion. For example, if $c_{k,l}=O(\epsilon)$ in (8.5), then $c_{N}$ in (F.3) and $c$ in (F.4) are $O(\epsilon)$.
If $D(\lambda=0,\theta,\epsilon)\equiv 0$, then we integrate by parts in (8.6) three times instead of two. The analog of (F.1) becomes
@@
(F.6) $|J(r)|\leq c\int_{0}^{\infty}\frac{\lambda^{2}}{1+|r\lambda|^{4}}\text{d}\lambda=O(r^{-3}),\ r\to\infty.$
@@
(F.8) $\begin{split}&\Upsilon(+\infty)+\Upsilon(-\infty)\\ &=\frac{1}{|\Phi_{x}^{\prime}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{\mathbb{R}}\text{sgn}(t-u)\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}u=0.\end{split}$

### v2

@@ C.2. Extracting the leading local singularity of $f_{\epsilon}$
(E.39) $\begin{split}&|\mathcal{Q}_{\epsilon}(u)|\leq c\epsilon^{1/2}(1+u^{2})^{-1},\ u\in\mathbb{R}.\end{split}$
For large $|u|$, (E.39) is proven by integrating by parts in (E.38) and using the bound on $\bar{h}_{\epsilon}$. For $|u|$ in a bounded set, we use the bound on $\bar{h}_{\epsilon}^{\prime\prime}$.
Set $\bar{h}_{\epsilon}(u)=h(u;\theta,\epsilon)-\mu h_{0}(u)$. Using (E.36) – (E.39), we get from (E.30)
(E.40) $\begin{split}&|G(\check{t},\check{q};\theta,\epsilon)-\mu G_{0}(\check{t},\check{q})|\leq c|J_{\epsilon}|+O(\epsilon^{1/2}),\\ &J_{\epsilon}:=\textstyle\sum_{j_{2}\in\mathbb{Z}}\mathcal{Q}_{\epsilon}(\check{t}+\check{q}-j_{2})(j_{2}-\check{q})_{+}^{1/2}=O(\epsilon^{1/2}).\end{split}$
@@
We evaluate the integrals in (7.19) similarly to [28, eqs. (3.16)–(3.20)]. Begin by considering
@@
Substitute $h_{0}$ expressed in terms of its Fourier transform, $\tilde{h}_{0}=\mathcal{F}h$, and use the identities [17, p. 360, eqs. 19 and 21]:
(E.42) $\begin{split}\mathcal{F}(x_{+}^{a-1})&=e^{ai\pi/2}\Gamma(a)(\lambda+i0)^{-a},\ a\not=0,-1,-2,\dots;\\ \mathcal{F}(1/x)&=i\pi\text{sgn}\,\lambda,\end{split}$
@@
Substitution of (E.44) back into (7.19) yields (7.20).
@@
We may assume $n=0$ in (8.5). The case $n\geq 1$ is completely analogous.
If 00 is not in the closure of $I$, integration by parts in (8.6) twice with respect to $\lambda$ yields the top line in (8.7). Therefore, using a partition of unity, we can assume that $I=(-\delta,\delta)$ for some small $0<\delta\ll 1$, and $D(\lambda,\theta,\epsilon)\in C_{0}^{\infty}(I)$ as a function of $\theta$. Change variable $\theta\to t=\sin\theta$ on $I$ to get
@@
and $\tilde{D}_{1}(\lambda,\mu,\epsilon)$ is the Fourier transform of $D_{1}(\lambda,t,\epsilon)$ with respect to $t$. Thus $\tilde{D}_{1}(\lambda,\mu,\epsilon)$ decays rapidly as $\mu\to\infty$, i.e. for any $N\in\mathbb{N}_{0}$ there exists $c_{N}$ such that
@@
A change of variable $\lambda\to r\lambda$ yields the top case in (8.7).
If $D(\lambda=0,\theta,\epsilon)\equiv 0$, then, away from $\theta=0$, we integrate by parts in (8.6) three times instead of two to prove the bottom line in (8.7). In a neighborhood of $\theta=0$, the analog of (F.1) becomes
@@
(F.6) $|J(r)|\leq c\int_{0}^{\infty}\frac{\lambda^{2}}{1+|r\lambda|^{4}}\text{d}\lambda=O(r^{-3}),\ r\to\infty,$
and the proof is complete.
@@
(F.8) $\begin{split}&\Upsilon(+\infty)+\Upsilon(-\infty)\\ &=\frac{1}{|\text{d}_{x}\Phi|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{\mathbb{R}}\text{sgn}(t-u)\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}\left(\frac{t}{|\text{d}_{x}\Phi|}\right)\text{d}t\text{d}u=0.\end{split}$